A population consists of the four numbers: 2, 3, 6, and 9. You have to consider all possible samples of size 2 that can be drawn from this population. Find the following:

a. the mean of the population;

b. the standard deviation of the mean of the population;

c. the mean of the sampling distribution of the sample means; and

d. the standard deviation of the sampling distribution of the sample means.


1
Expert's answer
2021-03-11T06:23:43-0500


(a),(b). Compute the mean and the standard deviation of the population. 

"E[X]=(2+3+6+9)\/4=5"

"E[X^2 ]=(2^2+3^2+6^2+9^2)\/4=32.5"

"Var[X]=E[X^2 ]-E[X]^2=32.5-25=7.5"

"\sigma(X)=\sqrt{Var[X] }=\sqrt{7.5}=2.7386"


Now we list all samples of size 2 and compute the mean for each sample. 



The sampling distribution of the sample means:

"P(\bar X_2=2.0)= P(\bar X_2=3.0)=P(\bar X_2=9.0)=1/16"

"P(\bar X_2=2.5)=P(\bar X_2=4.0)=P(\bar X_2=4.5)=P(\bar X_2=5.5)=P(\bar X_2=7.5)=1/8"

"P(\bar X_2=6/0)=3/16"


(c). Calculate the mean of the sampling distribution of the sample means. Compare this to mean of the population. 

"E[\bar X_2]=(2.0+3.0+9.0+2.5\cdot 2+4.0\cdot 2+4.5\cdot 2+5.5\cdot 2+7.5\cdot 2+6.0\cdot 3)/16=80/16=5.0=E[X]"


(d). Calculate the standard deviation of the sampling distribution of the sample means.

"E[\bar X_2^2] =(2.0^2+3.0^2+9.0^2+2.5^2\cdot 2+4.0^2\cdot 2+4.5^2\cdot 2+5.5^2\cdot 2+7.5^2\cdot 2+6.0^2\cdot 3)/16= 460/16=28.75"

"Var[\bar X_2 ]=E[\bar X_2^2 ]-E[\bar X_2 ]^2=28.75-5^2=3.75"

"\sigma(\bar X_2 )=\sqrt{Var[\bar X_2 ]}=\sqrt{3.75}=1.9365"